# Assimilative Causal Inference

### v1

@@
Assimilative Causal Inference
Marios Andreou Department of Mathematics, University of Wisconsin–Madison, 480 Lincoln Drive, Madison, WI 53706, USA Nan Chen [^1] Corresponding Author (chennan@math.wisc.edu) Department of Mathematics, University of Wisconsin–Madison, 480 Lincoln Drive, Madison, WI 53706, USA Erik Bollt Department of Mathematics and Department of Electrical and Computer Engineering, Clarkson University, 8 Clarkson Ave, Potsdam, 13699-5815, NY, USA
Summary
Causal inference determines cause-and-effect relationships between variables and has broad applications across disciplines. Traditional time-series methods often reveal causal links only in a time-averaged sense, while ensemble-based information transfer approaches detect the time evolution of short-term causal relationships but are typically limited to low-dimensional systems. In this paper, a new causal inference framework, called assimilative causal inference (ACI), is developed. Fundamentally different from the state-of-the-art methods, ACI uses a dynamical system and a single realization of a subset of the state variables to identify instantaneous causal relationships and the dynamic evolution of the associated causal influence range (CIR). Instead of quantifying how causes influence effects as done traditionally, ACI solves an inverse problem via Bayesian data assimilation, thus tracing causes backward from observed effects with an implicit Bayesian hypothesis. Causality is determined by assessing whether incorporating the information of the effect variables reduces the uncertainty in recovering the potential cause variables. ACI has several desirable features. First, it captures the dynamic interplay of variables, where their roles as causes and effects can shift repeatedly over time. Second, a mathematically justified objective criterion determines the CIR without empirical thresholds. Third, ACI is scalable to high-dimensional problems by leveraging computationally efficient Bayesian data assimilation techniques. Finally, ACI applies to short time series and incomplete datasets. Notably, ACI does not require observations of candidate causes, which is a key advantage since potential drivers are often unknown or unmeasured. The effectiveness of ACI is demonstrated by complex dynamical systems showcasing intermittency and extreme events.
Significance Statement
Causal inference is fundamental across scientific disciplines, yet existing methods struggle to capture instantaneous, time-evolving causal relationships in complex, high-dimensional systems. In this paper, assimilative causal inference (ACI) is developed, which is a paradigm-shifting framework that leverages Bayesian data assimilation to trace causes backward from observed effects. ACI solves the inverse problem rather than quantifying forward influence. It uniquely identifies dynamic causal interactions without requiring observations of candidate causes, accommodates short datasets, and scales efficiently to high dimensions. Crucially, it provides online tracking of causal roles, which may reverse intermittently, and facilitates a mathematically rigorous criterion for the causal influence range, revealing how far effects propagate. ACI opens new avenues for studying complex systems, where transient causal structures are critical.
@@ 1 Introduction
Causal inference determines cause-and-effect relationships between variables [28, 27, 36]. It has found wide applications across different disciplines such as atmospheric and ocean science, economics, and neuroscience [2, 33, 29]. In addition to discovering the interactions between variables, causal inference plays a significant role in model identification, policy evaluation, and decision-making [31, 5, 37, 49, 34].
@@
Methods for causal inference can be classified into different categories depending on the available resources for use. On the one hand, a natural way to identify the overall causal structures is to exploit multivariate time series, where temporal dependencies are often utilized to infer causal relationships. Methods falling into this category include Granger causality [16], transfer entropy [41, 3], mutual information [11, 15], convergent cross mapping [43, 26], causation entropy [45, 44], and mixture causal discovery [46]. On the other hand, models built upon physics can assist in understanding causal dependence between different variables. Information transfer based on the ensemble forecast of the underlying model has been exploited to indicate certain causal relationships for a short term [22]. In addition, linear response theory, which infers causal links by analyzing system responses to small perturbations at the equilibrium, has been utilized to reveal the attributions of variations in climate systems [23, 13, 18]. Other recently developed methods include using Koopman operators for causal discovery [38], causal graphs [20], and approaches based on machine learning [4, 30, 42].
@@
Nature can be regarded as a complex dynamical system where only one single random realization from this system is available. Due to the underlying strong nonlinearity and multiscale features, such a realization is usually intermittent and stochastic. Detecting instantaneous causal relationships as a function of time is crucial for understanding the dynamic interplay among variables, where the roles of causes and effects can shift repeatedly over time and at irregular intervals. In environmental science, extreme events, such as the triggering mechanisms of hurricanes and the abrupt transitions in weather patterns, have significant scientific and social impacts. Understanding the precursor of each extreme event and its subsequent implications requires real-time tracking of the causal relationship between them. Likewise, instantaneous causal analysis is of broad interest in neuroscience to discover temporary role reversals between brain regions during decision-making or cognitive tasks. However, most purely data-driven causal inference methods exploit data points over a long time series to reveal the average causal directions. In contrast, purely model-based methods, such as computing the information transfer in [22], utilize ensemble forecasts to infer the causal relationship when the system begins from a given initial condition and evolves towards statistical equilibrium. Yet, these methods are typically only able to provide time-dependent results for a very short time and become computationally challenging as the dimension of the system increases. Additionally, these model-based methods often struggle to optimally handle cases with a single observed realization.
@@
Given the importance of discovering instantaneous causal relationships, this paper introduces a new framework called assimilative causal inference (ACI) to compute time-dependent causal interactions in complex systems. As in many practical applications, only a single realization of a subset of the state variables is available as observations. Meanwhile, an associated model, often turbulent and stochastic, is accessible as a supplement. Fundamentally different from traditional causal inference methods, which treat causality as a forward problem (analyzing how causes propagate into effects), ACI addresses it as an inverse problem. Using Bayesian data assimilation, ACI traces causes backward from observed effects by quantifying whether incorporating information about the effects reduces uncertainty in reconstructing potential causes.
@@
ACI has several desirable advantages. First, it captures the evolving interplay among state variables, as their causal and effect roles can be reversed across time. Thus, ACI provides direct insights beyond time-averaged causal links, making it particularly valuable for studying turbulent systems with intermittency and regime switching. Second, ACI identifies the causal inference range (CIR), revealing how far causal effects propagate at each time instant. The framework includes a mathematically justified criterion to determine the CIR without empirical thresholds. While the importance of the CIR has been highlighted in complex networks [35, 50], ACI offers a unique and general method for its computation in turbulent systems with varying autocorrelation decay rates. Third, ACI scales efficiently to high-dimensional problems via Bayesian data assimilation, which is fundamentally different from information-transfer methods that are limited to low dimensions [22]. In addition, ACI can accommodate short time series and incomplete datasets, which often appear in geophysics and climate science. Notably, ACI does not require observations of candidate causes, which is a key advantage since potential drivers are usually unknown or unmeasured. Instead, uncertainty reduction in potential causal variables can be determined solely from the observed effects and governing dynamical model via Bayesian assimilation. As a final remark, while data assimilation has been used to assist in detecting the attribution of weather and climate-related events [17, 7], these studies primarily focus on estimating the state in response to specific external perturbations, rather than real-time causal inference or CIR estimation. ACI also fundamentally differs from these methods in its integration of data assimilation into causal inference.
@@
The remainder of this paper is organized as follows. Section 2 presents the ACI framework, followed by the discussion of CIR in Section 3. Section 4 extends the ACI framework to account for the presence of non-target variables. Numerical experiments on various nonlinear complex dynamical systems with intermittency and extreme events are presented in Section 5. Concluding remarks are contained in Section 6. Additional mathematical derivations and supplementary test results are available in the Supporting Information.
@@ 2.1 Setup and Notation
Consider a time interval $[0,T]$ and a specific time instant $t$, where $0\leq t\leq T$. Let $\mathbf{x}_{0\leq s\leq T}$ and $\mathbf{y}_{0\leq s\leq T}$ be two multivariate stochastic processes defined over this interval. Denote by $\mathbf{x}(t)$ and $\mathbf{y}(t)$ the corresponding multivariate random variables at time $t$. For notational simplicity, the dependence on $t$ is sometimes omitted when referring to these state variables and no distinction between random variables and their realizations is made. Let $\mathbf{x}(s\leq t)$ represent one realization of the stochastic process $\mathbf{x}_{0\leq s\leq T}$ as a time series over the interval $[0,t]$. Similarly, $\mathbf{x}(s\leq T)$ denotes a realization of $\mathbf{x}_{0\leq s\leq T}$ as a time series over the entire interval $[0,T]$. For simplicity, the time series are assumed to be continuously observed. An analogous framework can be developed for discrete-in-time observations. Throughout this paper, the time series is assumed to be free of observational noise.
According to Granger’s predictive causality or its nonlinear extension, transfer entropy, $\mathbf{y}$ is the cause of $\mathbf{x}$ if the knowledge of $\mathbf{y}(s\leq t)$ improves the prediction or reduces the uncertainty of $\mathbf{x}(t)$ [16, 41, 3]. Using information theory, this can be expressed as:
@@
where $\mathcal{S}(\cdot)$ is Shannon’s entropy and the history of the target variable in the conditioning, i.e., $\mathbf{x}(s<t)$, is omitted in (2.1) for notation simplicity.
@@ 2.2 Causal Discovery from a Bayesian Inverse Problem Viewpoint
A notable feature of (2.3) is that it establishes a connection between causal discovery and Bayesian data assimilation. Suppose the underlying turbulent and potentially stochastic model governing $\mathbf{x}$ and $\mathbf{y}$ is known. Running the model forward provides a statistical estimation of the state of $\mathbf{y}(t)$ (known as ‘forecast’). However, this state estimation differs when also incorporating the knowledge of the time series $\mathbf{x}(s\leq t)$ or $\mathbf{x}(s\leq T)$ (known as ‘analysis’), as the specific observation provides additional information to reduce the uncertainty when inferring the state of $\mathbf{y}(t)$. This two-step (forecast-analysis) process is known as Bayesian data assimilation [21, 32], where the distribution derived solely from the model statistics is referred to as the prior, while the observed time series is used to compute the likelihood, accounting for the uncertainty in the observational process. The two conditional distributions in (2.3) result from combining the model-based information with the observation-induced likelihood and are referred to as the posteriors. Specifically, the distribution on the left-hand side of (2.3) corresponds to smoothing, while the one on the right-hand side corresponds to filtering [12, 40].
@@ 2.3 ACI
Denote by $p_{t}^{\text{s}}$ and $p_{t}^{\text{f}}$ the smoother and filter posterior distributions at time $t$ (i.e., the two distributions on the left- and right-hand sides of (2.3)), respectively. If the uncertainty reduction in $p_{t}^{\text{s}}$ is more significant than that in $p_{t}^{\text{f}}$, then it is due to the incorporation of the future information of $\mathbf{x}$, which indicates towards the contribution from $\mathbf{y}(t)$ to the future states of $\mathbf{x}$.
Although the entropy difference has been widely used in many other causal inference methods, the relative entropy is employed hereafter to quantify the uncertainty reduction in $p_{t}^{\text{s}}$ related to $p_{t}^{\text{f}}$. In addition to the covariance (and other higher-order moments), which measure the variability of the random variable, the relative entropy also accounts for the difference in the mean state of the two distributions. This difference is crucial for reflecting the additional information gained by incorporating future observations. Furthermore, it is coordinate-free, i.e., it is invariant under general nonlinear changes of the state variables. Notably, relative entropy has been widely used to assess uncertainty reduction in the context of data assimilation [48, 24, 14, 9].
@@
Panel (a) of Figure 2.1 provides a high-level overview of the ACI framework, while Panel (b) presents a schematic illustration of ACI using the filter and smoother estimates. The Supporting Information contains more technical details.
Figure 2.1: The assimilative causal inference (ACI) framework. Panel (a): A high-level overview of the method. Panel (b): Schematic illustration of the ACI from a more technical viewpoint. Panel (c): ACI in the presence of non-target variables. Panel (d): Objective CIR, as an integration of the associated subjective CIRs.
@@ 3 Dynamic Discovery of the Causal Influence Range (CIR)
While the uncertainty reduction in the smoother solution related to the filter solution in (2.4) indicates an instantaneous causal link from $\mathbf{y}$ to $\mathbf{x}$ at time $t$, it does not reveal the temporal extent of this causal influence, namely, how many future units of $\mathbf{x}$ are affected by $\mathbf{y}(t)$. Since the memory of turbulent systems decays over time, $\mathbf{y}(t)$ is expected to effectively influence the future states $\mathbf{x}$ only within a limited time period, where this causal influence range (CIR) can vary significantly as a function of $t$. This raises two key questions: How can we determine the CIR and, more importantly, how can we do so objectively, i.e., without relying on empirical cutoff thresholds? This section outlines the core ideas for addressing these questions. Rigorous mathematical analysis and detailed algorithms are provided in the Supporting Information.

### v2

@@
Abstract
Causal inference is fundamental across scientific disciplines, yet existing methods struggle to capture instantaneous, time-evolving causal relationships in complex, high-dimensional systems. In this paper, assimilative causal inference (ACI) is developed, which is a methodological framework that leverages Bayesian data assimilation to trace causes backward from observed effects. ACI solves the inverse problem rather than quantifying forward influence. It uniquely identifies dynamic causal interactions without requiring observations of candidate causes, accommodates short datasets, and, in principle, can be implemented in high-dimensional settings by employing efficient data assimilation algorithms. Crucially, it provides online tracking of causal roles that may reverse intermittently and facilitates a mathematically rigorous criterion for the causal influence range, revealing how far effects propagate. The effectiveness of ACI is demonstrated by complex dynamical systems showcasing intermittency and extreme events. ACI opens valuable pathways for studying complex systems, where transient causal structures are critical.
@@ 1 Introduction
Causal inference determines cause-and-effect relationships between variables [pearl2000models, morgan2015counterfactuals, runge2019detecting]. It has found wide applications across different disciplines such as atmospheric and ocean science, economics, and neuroscience [angrist2009mostly, rubin1974estimating, peters2017elements]. In addition to discovering the interactions between variables, causal inference plays a significant role in model identification, policy evaluation, and decision-making [radebach2013disentangling, cafaro2015causation, runge2015identifying, yang2023finding, runge2019inferring].
@@
Methods for causal inference can be classified into different categories depending on the available resources for use. On the one hand, a natural way to identify the overall causal structures is to exploit multivariate time series, where temporal dependencies are often utilized to infer causal relationships. Methods falling into this category include Granger causality [granger1969investigating], transfer entropy [schreiber2000measuring, barnett2009granger], mutual information [cover1999elements, frenzel2007partial], convergent cross mapping [sugihara2012detecting, monster2017causal], causation entropy [sun2015causal, sun2014causation], and mixture causal discovery [varambally2023discovering]. On the other hand, models built upon physics can assist in understanding causal dependence between different variables. Information transfer based on the ensemble forecast of the underlying model has been exploited to indicate certain causal relationships for a short term [liang2005information]. In addition, linear response theory, which infers causal links by analyzing system responses to small perturbations at the equilibrium, has been utilized to reveal the attributions of variations in climate systems [lucarini2024detecting, falasca2023causal, hannart2016causal]. Other recently developed methods include using Koopman operators for causal discovery [rupe2024causal], causal graphs [kocaoglu2017experimental], dynamic neural relational inference [graber2020dynamic], and approaches based on machine learning [brand2023recent, prosperi2020causal, shanmugam2015learning].
@@
Nature can be regarded as a complex dynamical system where only a single random realization from this system is available. Due to the underlying strong nonlinearity and multiscale features, such a realization is usually intermittent and stochastic. Detecting instantaneous causal relationships as a function of time is crucial for understanding the dynamic interplay among variables, where the roles of causes and effects can shift repeatedly over time and at irregular intervals. In environmental science, extreme events, such as the triggering mechanisms of hurricanes and the abrupt transitions in weather patterns, have significant scientific and social impacts. Understanding the precursor of each extreme event and its subsequent implications requires online tracking of the causal relationship between them. Likewise, instantaneous causal analysis is of broad interest in neuroscience to discover temporary role reversals between brain regions during decision-making or cognitive tasks. However, most purely data-driven causal inference methods exploit data points over a long time series to reveal the average causal directions. In contrast, purely model-based methods, such as computing the information transfer in [liang2005information], utilize ensemble forecasts to infer the causal relationship when the system begins from a given initial condition and evolves toward statistical equilibrium. Yet, these methods are typically only able to provide time-dependent results for a very short time and become computationally challenging as the dimension of the system increases. Additionally, these model-based methods often struggle to optimally handle cases with a single observed realization.
@@
Given the importance of discovering instantaneous causal relationships, this paper introduces a methodological framework called assimilative causal inference (ACI) to compute time-dependent causal interactions in complex systems. As in many practical applications, only a single realization of a subset of the state variables is available as observations. Meanwhile, an associated model, often turbulent and stochastic, is accessible as a supplement. Fundamentally different from traditional causal inference methods, which treat causality as a forward problem (analyzing how causes propagate into effects), ACI addresses it as an inverse problem. Using Bayesian data assimilation, ACI traces causes backward from observed effects by quantifying whether incorporating information about the effects reduces uncertainty in reconstructing potential causes.
@@
ACI has several desirable advantages. First, it captures the evolving interplay among state variables, as their causal and effect roles can be reversed across time. Thus, ACI provides direct insights beyond time-averaged causal links, making it particularly valuable for studying turbulent systems with intermittency and regime switching. Second, ACI identifies the causal inference range (CIR), revealing how far causal effects propagate at each time instant. The framework includes a mathematically justified criterion to determine the CIR without empirical thresholds. While the importance of the CIR has been highlighted in complex networks [runge2012quantifying, ye2015distinguishing], ACI offers a unique and general method for its computation in turbulent systems with varying autocorrelation decay rates. Third, ACI scales efficiently to high-dimensional problems by leveraging computational methods developed in Bayesian data assimilation. This is fundamentally different from information-transfer methods limited to low dimensions [liang2005information]. In addition, ACI can accommodate short time series and incomplete datasets, which often appear in geophysics and climate science. Notably, ACI does not require observations of candidate causes, which is a key advantage since potential drivers are usually unknown or unmeasured. Instead, uncertainty reduction in potential causal variables can be determined solely from the observed effects and governing dynamical model via Bayesian assimilation. As a final remark, while data assimilation has been used to assist in detecting the attribution of weather and climate-related events [hannart2016dada, carrassi2017estimating], these studies primarily focus on estimating the state in response to specific external perturbations, rather than online causal inference or CIR estimation. ACI also fundamentally differs from these methods in its integration of data assimilation into causal inference.
@@
The remainder of this paper is organized as follows. Section 2 presents the operational details of the ACI framework, followed by the discussion of CIR in Section 3. Section 4 extends the ACI framework to account for the presence of non-target variables. Numerical results from applying ACI on various nonlinear complex dynamical systems with intermittency and extreme events are presented in Section 5. Concluding remarks are contained in Section 6. Additional mathematical derivations and supplementary test results are available in the Supplementary Information.
@@ 2 The Assimilative Causal inference (ACI) Framework
In this section, we provide a high-level overview of the methodology underlying the ACI framework, including its formulation of causal inference as a Bayesian inverse problem. Rigorous mathematical justifications are provided in the Supplementary Information.
@@ 2.1 Setup and Notation
Consider a time interval $[0,T]$ and a specific time instant $t$, where $0\leq t\leq T$. Let $\{\mathbf{x}(t)\}_{0\leq t\leq T}$ and $\{\mathbf{y}(t)\}_{0\leq t\leq T}$ be two multivariate stochastic processes defined over this interval. Denote by $\mathbf{x}(t)$ and $\mathbf{y}(t)$ the corresponding multivariate random variables at time $t$. Their evolution in time is defined by the following stochastic dynamical system:
$\frac{\mathrm{d}\mathbf{x}}{\mathrm{d}t}$ $=\mathbf{f}^{\mathbf{x}}(t,\mathbf{x},\mathbf{y})+\boldsymbol{\Sigma}^{\mathbf{x}}_{1}(t,\mathbf{x})\dot{\mathbf{W}}_{1}+\boldsymbol{\Sigma}^{\mathbf{x}}_{2}(t,\mathbf{x})\dot{\mathbf{W}}_{2},$ (2.1)
$\frac{\mathrm{d}\mathbf{y}}{\mathrm{d}t}$ $=\mathbf{f}^{\mathbf{y}}(t,\mathbf{x},\mathbf{y})+\boldsymbol{\Sigma}^{\mathbf{y}}_{1}(t,\mathbf{x},\mathbf{y})\dot{\mathbf{W}}_{1}+\boldsymbol{\Sigma}^{\mathbf{y}}_{2}(t,\mathbf{x},\mathbf{y})\dot{\mathbf{W}}_{2},$
where $\dot{\mathbf{W}}_{1}$ and $\dot{\mathbf{W}}_{2}$ are independent mean-zero Gaussian random vectors with uncorrelated components that possess unit variance. Mathematical details regarding the structure of (2.1) are given in the Supplementary Information. For notational simplicity, the dependence on $t$ is sometimes omitted when referring to these state variables and no distinction between random variables and their realizations is made. Let $\mathbf{x}(s\leq t)$ represent one realization of the stochastic process $\{\mathbf{x}(t)\}_{0\leq t\leq T}$ as a time series over the interval $[0,t]$. Similarly, $\mathbf{x}(s\leq T)$ denotes a realization of $\{\mathbf{x}(t)\}_{0\leq t\leq T}$ as a time series over the entire interval $[0,T]$. For simplicity, the time series are assumed to be continuously observed. An analogous framework can be developed for discrete-in-time observations. Throughout this paper, the time series is assumed to be free of observational noise.
According to Granger’s predictive causality or its nonlinear extension, transfer entropy, $\mathbf{y}$ is the cause of $\mathbf{x}$ at time $t$ if the knowledge of $\mathbf{y}(s\leq t)$ improves the prediction or reduces the uncertainty of $\mathbf{x}(t)$ [granger1969investigating, schreiber2000measuring, barnett2009granger]. Using information theory, this can be expressed as:
@@
where $\mathcal{S}(\cdot)$ is Shannon’s entropy and the history of the target variable in the conditioning, i.e., $\mathbf{x}(s<t)$, is omitted in (2.2) for notational simplicity.
@@ 2.2 Causal Discovery from a Bayesian Inverse Problem Viewpoint
A notable feature of (2.4) is that it establishes a connection between causal discovery and Bayesian data assimilation. Suppose the underlying turbulent and potentially stochastic model governing $\mathbf{x}$ and $\mathbf{y}$ is known and given by (2.1). Running the model forward provides a statistical estimation of the state of $\mathbf{y}(t)$ (known as ‘forecast’). However, this state estimation differs when also incorporating the knowledge of the time series $\mathbf{x}(s\leq t)$ or $\mathbf{x}(s\leq T)$ (known as ‘analysis’), as the specific observation provides additional information to reduce the uncertainty when inferring the state of $\mathbf{y}(t)$. This two-step (forecast-analysis) process is known as Bayesian data assimilation [law2015data, reich2015probabilistic], where the distribution derived solely from the model statistics is referred to as the prior, while the observed time series is used to compute the likelihood, accounting for the uncertainty in the observational process. The two conditional distributions in (2.4) result from combining the model-based information with the observation-induced likelihood and are referred to as the posteriors. Specifically, the distribution on the left-hand side of (2.4) corresponds to smoothing, while the one on the right-hand side corresponds to filtering [evensen2022data, sarkka2023bayesian].
@@ 2.3 ACI
Denote by $p_{t}^{\text{s}}$ and $p_{t}^{\text{f}}$ the smoother and filter posterior distributions at time $t$ (i.e., the two distributions on the left- and right-hand sides of (2.4)), respectively. If the uncertainty reduction in $p_{t}^{\text{s}}$ is more significant than that in $p_{t}^{\text{f}}$, then it is due to the incorporation of the future information of $\mathbf{x}$, which indicates toward the contribution from $\mathbf{y}(t)$ to the future states of $\mathbf{x}$.
Although the entropy difference has been widely used in many other causal inference methods, the relative entropy is employed hereafter to quantify the uncertainty reduction in $p_{t}^{\text{s}}$ related to $p_{t}^{\text{f}}$. In addition to the covariance (and other higher-order moments), which measure the variability of the random variable, the relative entropy also accounts for the difference in the mean state of the two distributions. This difference is crucial for reflecting the additional information gained by incorporating future observations. Furthermore, it is coordinate-free, i.e., it is invariant under general nonlinear changes of the state variables. Notably, relative entropy has been widely used to assess uncertainty reduction in the context of data assimilation [xu2007measuring, majda2018model, fowler2013observation, chen2023stochastic].
@@
Under (2.5), ACI provides a unique formulation for identifying causal relationships. In traditional approaches, causal links are inferred by running (2.1) forward to quantify the informational response on the target variables’ entropy when conditioned on the history of the candidate causes. ACI instead solves an inverse problem, thus tracing causes back from the data of the observed subsequent effects. By interpolating causality from effects to causes instead of extrapolating the latter forward-in-time, ACI offers an inherently more robust framework compared to classical predictive causality. Interpolation is typically more stable because it operates within the range of observed information, whereas extrapolation requires projecting beyond available data and is therefore far more sensitive to model error and uncertainty.
Panel (a) of Figure 1 provides a high-level overview of the ACI framework, while Panel (b) presents a schematic illustration of ACI using the filter and smoother estimates. The Supplementary Information contains more technical details.
Figure 1: The ACI framework. Panel (a): A high-level overview of the proposed method. Panel (b): Schematic illustration of ACI from a more technical viewpoint. Panel (c): ACI in the presence of non-target variables. Panel (d): Objective CIR, as an average of the associated subjective CIRs.
@@ 3 Dynamic Discovery of the Causal Influence Range (CIR)
While the uncertainty reduction in the smoother solution related to the filter solution in (2.5) indicates an instantaneous causal link from $\mathbf{y}$ to $\mathbf{x}$ at time $t$, it does not reveal the temporal extent of this causal influence, namely, how many future units of $\mathbf{x}$ are affected by $\mathbf{y}(t)$. Since the memory of turbulent systems decays over time, $\mathbf{y}(t)$ is expected to effectively influence the future states $\mathbf{x}$ only within a limited time period, where this causal influence range can vary significantly as a function of $t$. This raises two key questions: How can we determine the CIR and, more importantly, how can we do so objectively, i.e., without relying on empirical cutoff thresholds? This section outlines the core ideas for addressing these questions. Rigorous mathematical analysis and detailed algorithms are provided in the Supplementary Information.